2.76 g of silver carbonate on being strongly heated yields a residue weighing
a. 2.16 g
b. 2.48 g
c. 2.32 g
d. 2.64 g

Grade:10

6 Answers

Jitender Pal
9 years ago
Sol. Unlike other metal carbonates that usually decomposes into metal oxides liberating carbon dioxide, silver carbonate on heating decomposes into elemental silver liberating mixture of carbon dioxide and oxygen gas as Ag2CO3 (s) 2Ag (s) + CO2 (g) + 1/2 O2 (g) MW = 276g 2 × 108 = 216 g Hence, 2.76 g of Ag2CO3 on heating will give 216/276× 2.76 = 2.16 g Ag as residue.
Ayushman Singh
6 years ago
the rxn is as follows Ag2CO3=Ag + Co2 + O2 No need to balance the equation just apply POAC. 2×moles of Ag2Co3= 1× Mole of Ag and moles = Weight in gram/molr mass put the values and u will get the same answer. POAC is very simple just conserve the no of atoms of that particular element on both the sides and get the answer.
Janki
6 years ago
Here, First write the equation..Ag2CO3 --------》 2Ag + CO2 + 1/2 O2263.6g 203.6g (standard molecular wt)But, Ag2CO3=2.76gSo, 263.6g Ag2CO3 ------ 203.6g 2Ag 2.76g Ag2CO3. ----- ?So, 2.76 × 203.6/263.6 =2.13g ~2.16gHope it helped! :)
Mazid
5 years ago
First of all we break Ag2Co3 by decomposition reaction then we get Ag2Co3 -> 2Ag + Co2 + 1/2O2Now Molar mass of Ag is = 2x108 =216gHence 2.76g of Ag2Co3 give 216/276x2.76 = 2.16g of residue remain
Arun kumar
4 years ago
Ag2CO3---2Ag+CO2+1/2O2
276gm gives-----216 gm.  Then
2.76gm gives----2.16gm.  Ans.........................
Rishi Sharma
3 years ago
Hello students,
The solution is as follows
Unlike other metal carbonates that usually decomposes into metal oxides liberating carbon dioxide, silver carbonate on heating decomposes into elemental silver liberating mixture of carbon dioxide and oxygen gas as
Ag2CO3 (s) --> 2 Ag (s) + CO2 (g) + 1/2 O2 (g)
276g of Ag2CO3 on heating will give (2 × 108 =) 216 g
Hence, 2.76 g of Ag2CO3 on heating will give 216/276× 2.76 = 2.16 g Ag as residue.
Thanks
I hope the above solution will solve all your doubts.
All the best for your exams.